### v1

# Democratizing Large Language Models via Personalized Parameter-Efficient Fine-tuning

Abstract.
Personalization in large language models (LLMs) is increasingly important, aiming to align LLM’s interactions, content, and recommendations with individual user preferences. Recent advances in LLM personalization have spotlighted effective prompt design, by enriching user queries with non-parametric knowledge through behavior history retrieval and textual profiles. However, these approaches were limited due to a lack of model ownership, resulting in constrained customization and privacy issues. Moreover, they often failed to accurately capture user behavior patterns, especially in cases where user data were complex and dynamic. To address these shortcomings, we introduce One PEFT Per User (OPPU) [^1] The code is available at https://github.com/TamSiuhin/OPPU, which employs personalized parameter-efficient fine-tuning (PEFT) modules, to store user-specific behavior patterns and preferences. By plugging in users’ personal PEFT parameters, they can own and use their LLMs personally. OPPU integrates parametric user knowledge in the personal PEFT parameters with the non-parametric knowledge acquired through retrieval and profile. This integration adapts individual LLMs to user behavior shifts. Experimental results demonstrate that OPPU significantly outperforms existing prompt-based methods across seven diverse tasks in the LaMP benchmark. Further in-depth studies reveal OPPU’s enhanced capabilities in handling user behavior shifts, modeling users at different active levels, maintaining robustness across various user history formats, and displaying versatility with different PEFT methods.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Personalization refers to mining users’ behavior history, and therefore tailoring and customizing a system’s interactions, content, or recommendations to meet specific needs, preferences, and characteristics of individual users (Tan and Jiang, 2023; Chen et al., 2023; Chen, 2023). By adapting to each user’s preferences, personalization systems enhance user experience, increasingly getting vital in areas like content recommendation (Qian et al., 2013; Li et al., 2023a; Wu et al., 2023; Baek et al., 2023), user simulation (Dejescu et al., 2023; Zhang and Balog, 2020), personalized chatbots (Srivastava et al., 2020; Ma et al., 2021; Qian et al., 2013), user profiling (Gu et al., 2020; Gao et al., 2023), healthcare (Johnson et al., 2021; Goldenberg et al., 2021), and education (Alamri et al., 2021; Pratama et al., 2023).
Large language models (LLMs) display emergent abilities not seen in smaller models (Wei et al., 2022a; Lu et al., 2023), as they have billions of parameters and are trained on vast corpora. These abilities include step-by-step reasoning (Wei et al., 2022b), in-context learning (Min et al., 2022), and instruction following (Wei et al., 2021). However, existing LLMs predominantly follow the “one-size-fits-all” paradigm. They are generally trained on extensive, domain-agnostic datasets, which limits their effectiveness in meeting the specific needs and preferences of individual users (Chen et al., 2023; Salemi et al., 2023). Therefore, the challenge of integrating the strong generative capabilities of LLMs with the tailored requirements of individual users has emerged as a significant area of research (Jang et al., 2023; Li et al., 2023b).
Existing works on personalizing LLMs have predominantly concentrated on the development of prompt templates. These prompt-based personalization methods fall into three categories: vanilla personalized prompt, retrieval-augmented personalized prompt, and profile-augmented prompt. The vanilla personalized prompt approach leverages the in-context learning capability of LLMs, utilizing the user’s entire or random sampled behavior history as contextual examples (Zhiyuli et al., 2023; Dai et al., 2023; Christakopoulou et al., 2023). Considering the growing length of user behavior history and the limited context window of LLMs, some studies have applied retrieval methods to select the most relevant part of user behavior history to enhance LLM personalization (Salemi et al., 2023; Mysore et al., 2023). Besides the retrieval, some techniques explicitly generate user preferences and profiles in natural language to augment LLMs’ input. For instance, Richardson et al. (2023) proposed to employ instruction-tuned LLMs, e.g., ChatGPT, to summarize user preferences and behavior patterns based on their history content.
Figure 1. LLM ownership and behavior shift are two challenges that developing personalized LLMs has to face. Ownership emphasizes that the model needs to be owned by individual user to enhance customization and privacy. Behavior shift adaption refers to the LLMs’ ability to effectively generalize and adapt to emerging new patterns in user behaviors.
Despite much research progress has been made in LLM personalization, existing methods face ownership and behavior shift challenges, which are illustrated in Figure 1:
  • Ownership: Existing methods are processed in a centralized way, where user history is encoded in a personalized prompt and processed by centralized LLMs. This paradigm limits the model’s customization and ability to provide deep, personalized experiences tailored to individual users. Moreover, when using a centralized model, users often have to share personal data with the service provider, which raises concerns about how user data are stored, used, and protected.
  • Behavior Pattern Generalization: As is revealed by Shi et al. (2023), LLMs can be easily distracted by irrelevant context information that retrieval can hardly avoid. In the realm of LLM personalization, where the retrieval corpus was confined to a specific user’s behaviors, retrieval augmentation might underperform, especially when the user’s past behaviors did not closely mirror the patterns needed for the query at hand.
In light of these challenges, we propose One PEFT Per User (OPPU), where each user is equipped with a personalized, parameter-efficient fine-tuning (PEFT) module. Characterized by PEFT’s plug-and-play functionality and the minimal weight of updated parameters (typically less than 1% of the base LLM), OPPU facilitates LLM ownership and exhibits superior generalization in scenarios of user behavior shifts. By fine-tuning the PEFT module with the user’s personal behavior history, the personalized PEFT parameters encapsulate behavior patterns and preferences. This process, when integrated into base LLMs, allows users to obtain their private LLMs, ensuring LLM ownership and enhancing model customization. Furthermore, as is revealed by Gupta et al. (2024), fine-tuning LLMs is more effective than retrieval augmentation when the retrieved instances are not highly relevant to the query. The fine-tuned personal LLMs in OPPU are adept at capturing complex behavior patterns and thus capable of understanding new behaviors with less reliance on highly relevant history data.
Experimental results show that OPPU achieves state-of-the-art performance on all seven public tasks in the Language Model Personalization (LaMP) benchmark (Salemi et al., 2023). Additional studies highlight the importance of integrating non-parametric user knowledge, sourced from retrieved user history, with parametric user knowledge from personal PEFT parameters. Notably, in scenarios of user behavior shift, where the user history is less relevant to the current user query at hand, OPPU significantly outperforms retrieval-based methods. Moreover, OPPU exhibits strong resilience against varying user history formats and demonstrates versatility across different PEFT methods, among other advantages.
To summarize, the contribution of OPPU lies in its pioneering approach to PEFT-based LLM personalization. Each user (or user cohort) benefits from a personal PEFT module, which not only ensures LLM ownership but also significantly improves the model’s ability to adapt to shifts in user behavior. The superiority of OPPU is evidenced by state-of-the-art performance across seven tasks in the LaMP benchmark. By introducing this innovative parametric-based personalization technique, OPPU opens up new opportunities in the realm of democratizing personalized LLMs.
2. Related Work
2.1. Personalization of LLMs
The thrust of existing LLM personalization research is centered on designing prompt that incorporate historical user-generated content and behavior. These approaches help LLMs understand users’ preferences, tailoring responses to individual needs (Tan and Jiang, 2023; Chen et al., 2023). The endeavors towards personalized LLMs mainly fall into three categories: vanilla personalized prompts, retrieval-augmented personalized prompts, and profile-augmented personalized prompts.
In the vanilla personalized prompt category, researchers use in-context and few-shot learning to encode either complete or a sample of user behavior history as contextual examples. For instance, Dai et al. (2023) and Kang et al. (2023) encode the user’s personal rating history as few-shot demonstration examples. Liu et al. (2023b) supply LLMs with user’s interaction history related to specific tasks to help LLMs generate personalized content. PerSE (Wang et al., 2023) design prompt to do personalized story assessment by presenting a few exemplary reviews from the review. BookGPT (Zhiyuli et al., 2023) employs a few-shot prompt strategy to make LLMs understand the correlation between book content and personalized prompting. Moreover, some research works (Christakopoulou et al., 2023; Zhiyuli et al., 2023) also discovered a long user history would bring better performance. To manage the growing user behavior data and LLMs’ limited context windows, the retrieval-augmented personalized prompt approach has emerged. For instance, LaMP (Salemi et al., 2023) introduces a retrieval-augmented method to obtain the most relevant content in the user’s behavioral history and incorporate it into the prompt. AuthorPred (Li et al., 2023b) utilizes retrieve relevant past user-written documents for personalized text generation. Pearl (Mysore et al., 2023) proposes a generation-calibrated retriever to select historic user-authored documents for prompt augmentation. Moving beyond simple retrieval, some researchers summarize user preferences and behavior patterns into natural language profiles for input query augmentation, termed profile-augmented personalized prompts. Richardson et al. (2023) use the instruction-tuned LLMs to generate an abstract summary of user history data, augmenting retrieval-based personalization methods. ONCE (Liu et al., 2023a) employs LLMs to generate users’ topics and regions of interest based on their browsing history as user profiles, aiding LLMs in preference capture in downstream tasks. There is also another line of work focusing on designing personalized alignment methods via parameter merging (Jang et al., 2023) and personalized reward model (Cheng et al., 2023).
Previous works largely hinge on prompt design, integrating retrieval and user profiles. However, existing approaches are limited by model ownership and users’ behavior shifts. Our OPPU introduces personalization at the parametric level, via the personal PEFT module, which ensures model ownership and superior generalization, especially in cases with less relevant user history for retrieval.
2.2. Parameter-Efficient Fine-tuning (PEFT)
With the exponentially growing parameters in LLMs, fine-tuning all parameters is expensive (Liu et al., 2022a; Xu et al., 2023; Gupta et al., 2024). To mitigate this gap, a few lightweight alternatives, known as parameter-efficient fine-tuning (PEFT) have been proposed to update only a small number of extra parameters while keeping the pretrained weights frozen to save the computes (He et al., 2021; Fu et al., 2023). For example, adapter tuning (Houlsby et al., 2019) injects learnable parameters in the models’ each feedforward layer, and only the plug-in parameters are updated at the fine-tuning stage. Inspired by the success of discrete textual prompt (Sanh et al., 2022; Wang et al., 2022), prefix tuning (Li and Liang, 2021) and prompt tuning (Lester et al., 2021) are proposed to learn the optimized prompt and prefix for specific use via fine-tuning. LoRA (Hu et al., 2021) proposes to add low-rank matrices on pretrained weights to approximate parameter updates. $\rm{(IA)^{3}}$ (Liu et al., 2022b) plugs in learned vectors to scale the activation in the attention mechanism for efficient fine-tuning. These approaches achieve comparable performance to full parameter fine-tuning by updating less than 1% of the original LLM parameters. Besides parameter savings, PEFT methods are effective in combating catastrophic forgetting (Pfeiffer et al., 2021) and robust to out-of-distribution samples (Li and Liang, 2021).
The small number of updated parameters and plug-and-play nature of PEFT make it an ideal solution for efficient LLM personalization and constitute model ownership. Our work pioneers the concept of storing user history within personal PEFT parameters. Each user is equipped with a unique, easily integrable PEFT module, serving as a key to their own personalized LLMs.
Figure 2. Overview of our proposed OPPU, where each user is equipped with a personal PEFT module and plug-in base LLMs to get their individual LLM. Beyond parametric personalization via PEFT, OPPU is also compatible with the non-parametric user knowledge via retrieval and profile augmentation.
3. Proposed Approach
3.1. Research Problem Formulation
Generative language models generally take an input token sequence $x$ and output a sequence of tokens $y$ that follows the $x$. For the personalization of LLMs, happening at time $t$, the language model’s output $y_{u}$ specifically for user $u$ is conditioned on both input $x_{u}$ and the user $u$’s behavior history $\mathcal{H}_{u}$. To elaborate, $\mathcal{H}_{u}^{t}=\{h_{u}^{t_{i}}\},t_{i}<t$, where the user’s history data $\mathcal{H}_{u}$ contains all user behavior $h^{t_{i}}$ happened before query time $t$. More specifically, the user behavior $h_{u}^{t_{i}}$ may consist of $(x_{u}^{t_{i}},y_{u}^{t_{i}})$ pairs, mirroring the task-specific query-answer format $(x_{u},y_{u})$. Alternatively, $h_{u}^{t_{i}}$ could be text sequences that provide context for the user’s behavior patterns without necessarily conforming to the task’s format. Overall, each data entry contains three components: input sequence $x_{u}$ given by the user, user’s history behavior $\mathcal{H}_{u}$ that contains the user’s behavior that happened before the input query, and a target output $y_{u}$ specifically for user $u$ that the model is expected to produce.
3.2. Base LLMs Task Adaption
Given that off-the-shelf LLMs are not inherently equipped for personalization tasks, we align with the methods of LaMP (Salemi et al., 2023) and Richardson et al. (2023) by fine-tuning LLMs for fair comparison. In this section, we will describe the process of developing base LLMs, a critical step to enhance their general capability in comprehending and executing the specific personalization tasks required, without specific user’s preference or bias.
Non-personalized base LLM ($\mathrm{LLM_{NP}}$) ignores the user’s behavior history and serves as a baseline. It is fine-tuned on $\{(\phi_{p}(x_{u}),y_{u})\}$ data pairs, where $\phi_{p}$ is the prompt construction function. This approach only contains the task-related data and ignores the user behavior history, therefore making it non-personalized.
Retrieval-augmented base LLM ($\mathrm{LLM_{RAG}}$) is fine-tuned on input that includes the retrieved top-$k$ relevant user history items, which the input sequence $x^{\prime}$ can be defined as:
(1) $\overline{x_{u}}=\phi_{p}(x_{u},\mathcal{R}(x_{u},\mathcal{H}_{u},k)),$
where $\mathcal{R}$ denotes the retriever, $\phi_{p}$ denotes the prompt construction function. This approach aims to enhance personalization by integrating the most pertinent user history items directly into the LLM’s prompt.
Profile-Augmented Base LLM ($\mathrm{LLM_{PAG}}$) enhances its input by appending a natural language user profile that describes the user’s preferences and behavior patterns. This profile can be added to the non-personalized and retrieval-augmented prompts. The input $x^{\prime}$ to LLM could be denoted as:
(2) $\overline{x_{u}}=\phi_{p}(x_{u},\mathcal{R}(x_{u},\mathcal{H}_{u},k),s_{u}),$
where $s_{u}$ is a textual user profile that describes user’s preference and behavior patterns that are generated by an instruction-tuned LLM (Vicuna (Chiang et al., 2023) or ChatGPT (Ouyang et al., 2022)) based on the user history data.
To make this process more computationally efficient, we adopt the low-rank adaptation (LoRA) (Hu et al., 2021) for base modeling task adaption that only updates about 0.5% of external parameters compared to the total LLM parameter size. After training the LoRA, we merge the LoRA parameters to the base model to get base LLMs equipped with the capabilities for the corresponding tasks.
3.3. One PEFT Per User (OPPU)
After obtaining the base model that comprehends the responding task, in real-world deployment, users can only assess the parameter of the base model and their personal behavior history data, controlling the privacy risks. In this section, we will introduce how to make personalized LLMs via parameter-efficient fine-tuning (PEFT), and its integration with non-parametric personalization, i.e., retrieval-augmentation and profile-augmentation. To better understand the parametric personal knowledge in PEFT parameters and non-parametric personalized knowledge in the prompt, we decompose OPPU into parametric knowledge only and the integration of both parametric and non-parametric personalized knowledge.
3.3.1. Inject Parametric Personalized Knowledge via PEFT
We first explore injecting the parametric personalized knowledge into base LLMs and exclude the non-parametric augmentation method to focus purely on the impact of parametric modifications. In this scenario, the user history data are all stored in the PEFT module’s parameters thus enhancing privacy preservation. For each user $u$, OPPU would maintain a private PEFT module $\mathrm{PEFT}{u}$, tailored to capture and adapt to the user’s behavioral patterns as reflected in their historical data. More specifically, consider a user $u$ with history $\mathcal{H}_{u}$, the personalized PEFT module is plugged in base LLM ${LLM_{NP}}$ and optimized on the user’s personal history data. The input sequence $x^{\prime}$ for LLMs is $x_{u}^{\prime}=\phi_{p}(x^{t_{i}}_{u})$ and optimized using the standard cross entropy loss against the corresponding output $y^{t_{i}}_{u}$, where $(x^{t_{i}}_{u},y^{t_{i}}_{u})\in\mathcal{H}_{u}$. This procedure could capture the user behavior pattern presented in all user behavior history more comprehensively store it in plug-and-play personal PEFT parameters, constituting a personalized LLM owned by users.
3.3.2. Integrating Non-Parametric & Parametric Knowledge
Inspired by recent research on debating between fine-tuning versus retrieval-augmented generation (Gupta et al., 2024), as well as insights into the integration of non-parametric knowledge through retrieval and parametric knowledge via PEFT, we proceed to investigate the integration of our personalized PEFT parametric user knowledge with non-parametric user knowledge from both retrieval and profile augmentation methods.
Integrating with Retrieval Augmentation
Our first step involves integrating the personal PEFT module with retrieval augmentation, in which the $\mathrm{LLM_{RAG}}$ acts as the base LLM and each user’s plug-in PEFT parameters are updated on the following input sequence $x^{\prime}$:
(3) ${{x}_{u}^{t_{i}}}^{\prime}=\phi_{p}(x_{u}^{t_{i}},\mathcal{R}(x_{u}^{t_{i}},\mathcal{H}_{u}^{t_{i}},k)),$
where $\phi{p}$ denotes the prompt construction function and $\mathcal{R}$ represents the retriever. The retriever selects the top $k$ relevant history items from the user’s behavior history corpus $\mathcal{H}_{u}^{t_{i}}$ in response to the user history query $x_{u}$. It’s important to note that the retrieval is based on past behavior that occurred before the current query $x_{u}^{t_{i}}$. The PEFT parameters are updated using cross entropy loss against the corresponding historical user response $y_{u}^{t_{i}}$.
Integrating with User Profile Augmentation
Following Richardson et al. (2023), we also incorporate profile augmentation into our method for injecting non-parametric knowledge into prompts. In this scenario, the personal PEFT parameter $\mathrm{PEFT_{u}}$ is plugged in $\mathrm{LLM_{PAG}}$ and optimized using the input ${{x}_{u}^{t_{i}}}^{\prime}$:
(4) ${{x}_{u}^{t_{i}}}^{\prime}=\phi_{p}(x_{u}^{t_{i}},\mathcal{R}(x_{u}^{t_{i}},\mathcal{H}_{u}^{t_{i}},k),s_{u}),$
which includes the user’s profile $s_{u}$ on the basis of retrieval augmentation. This profile $s_{u}=LLM(\mathcal{H}_{u})$ is a summary of the user’s preferences and behavior patterns, generated by an LLM that can follow human instruction.
By integrating the user knowledge in both private PEFT module and non-parametric retrieval and profile methods, OPPU is better equipped to understand a user’s behavior patterns. This integrated approach enhances the model’s ability to generalize these patterns to new and emerging user behaviors.
3.3.3. Adapting to Behavior History Beyond Task Conformity
It is worth mentioning that the user behavior history does not always align neatly with the format of a user’s query. For instance, in personalized tweet paraphrasing tasks, where to input a sequence of text $x_{u}$, the desired output is the corresponding paraphrased text $y_{u}$ concerning user history $\mathcal{H}_{u}$. However, $\mathcal{H}_{u}$ often comprises only the user’s historical tweets $\{y_{u}^{t_{i}}\}\in\mathcal{H}_{u}$, not the $(x_{u}^{t_{i}},y_{u}^{t_{i}})$ pairs. In scenarios where user history does not directly provide preferences in the specific task format, we find tuning a personal PEFT could also provide LLMs with more context and benefit the personalization task performance. Specifically, we update the personal PEFT parameters using unsupervised pretraining loss to predict next tokens over the user history sequences $h_{u}^{t_{i}}\in\mathcal{H}_{u}$. This approach infuses personal historical knowledge into the LLM via the integrated PEFT parameters.
Overall, we envision the proposed OPPU as a versatile LLM personalization framework, where each user possesses their own PEFT parameters that contain their personal behavior history and preference. By plugging their personal PEFT parameters into the base LLMs, users could get their personalized LLMs, while achieving a better understanding of users’ personality from the parametric dimension.
4. Experiments
4.1. Experimental Settings
[…]
Table 2. Main experiment results on the LaMP benchmark. R-1 and R-L denote ROUGE-1 and ROUGE-L, respectively. $k$ refers to the number of retrieved items, with $k=0$ indicating no retrieval. $\uparrow$ indicates that higher values are better, and $\downarrow$ implies lower values are preferable. For each task, the best score is in bold and the second best is underlined.
Task Metric Non-Personalized RAG PAG RAG+OPPU (Ours) PAG+OPPU (Ours) k=0 Random k=1 k=2 k=4 k=0 k=1 k=0 k=1 k=2 k=4 k=0 k=1 LaMP-1: Personalized Citation Identification Accy $\uparrow$ 0.659 0.650 0.659 0.691 0.691 0.756 0.755 0.683 0.675 0.707 0.723 0.772 0.797 F1 $\uparrow$ 0.657 0.647 0.657 0.689 0.690 0.755 0.755 0.682 0.674 0.705 0.723 0.772 0.794 LaMP-2N: Personalized News Categorization Acc $\uparrow$ 0.787 0.785 0.820 0.832 0.832 0.817 0.817 0.810 0.823 0.834 0.838 0.827 0.831 F1 $\uparrow$ 0.538 0.527 0.598 0.632 0.647 0.623 0.621 0.589 0.615 0.635 0.661 0.648 0.638 LaMP-2M: Personalized Movie Tagging Acc $\uparrow$ 0.478 0.499 0.587 0.598 0.622 0.534 0.587 0.600 0.626 0.634 0.645 0.636 0.648 F1 $\uparrow$ 0.425 0.441 0.512 0.514 0.542 0.476 0.506 0.493 0.531 0.535 0.553 0.536 0.540 LaMP-3: Personalized Product Rating MAE $\downarrow$ 0.223 0.259 0.214 0.214 0.232 0.321 0.223 0.179 0.196 0.214 0.223 0.205 0.143 RMSE $\downarrow$ 0.491 0.590 0.535 0.463 0.535 0.582 0.473 0.443 0.518 0.463 0.526 0.473 0.378 LaMP-4: Personalized News Headline Gen. R-1 $\uparrow$ 0.186 0.187 0.191 0.196 0.198 0.187 0.193 0.904 0.194 0.196 0.199 0.189 0.194 R-L $\uparrow$ 0.167 0.168 0.172 0.176 0.178 0.168 0.173 0.171 0.175 0.177 0.180 0.170 0.175 LaMP-5: Personalized Scholarly Title Gen. R-1 $\uparrow$ 0.476 0.478 0.505 0.510 0.499 0.486 0.516 0.519 0.522 0.511 0.526 0.490 0.525 R-L $\uparrow$ 0.415 0.418 0.445 0.444 0.434 0.429 0.440 0.442 0.457 0.440 0.467 0.428 0.473 LaMP-7: Personalized Tweet Paraphrasing R-1 $\uparrow$ 0.527 0.524 0.568 0.577 0.562 0.542 0.568 0.539 0.579 0.575 0.581 0.542 0.577 R-L $\uparrow$ 0.474 0.474 0.521 0.527 0.514 0.501 0.518 0.483 0.533 0.531 0.528 0.492 0.533
4.1.1. Datasets
We adopt the Large Language Model Personalization (LaMP) benchmark (Salemi et al., 2023) for our experiments, which consists of seven public language model personalization tasks, including four classification tasks (personalized citation identification, personalized movie tagging, personalized producing rating, personalized news categorization) and three generation tasks (personalized news headline generation, personalized scholarly title generation, personalized tweet paraphrasing). [^2] We exclude the LaMP-6: personalized Email subject generation task since it involves private data that we don’t have access to. We mainly focus on the most active uses and select 100 users from the time-based dataset version that have the longest history log as the test set, while all other users are used for base LLM training. Statistics of datasets are in Table 1.
4.1.2. Baselines
We compare our proposed OPPU with the non-personalized baseline and the retrieval-augmented and profile-augmented LLM personalization methods.
  • Non-Personalized Baseline: We present two approaches under the non-personalized setting: non-retrieval and random history. Non-retrieval method refers to only feeding the user’s query without revealing the user’s behavior history to the LLMs. Random history baseline means augmenting the user’s query with random history behavior from all user history corpus.
  • Retrieval-Augmented Personalization (RAG): We follow the retrieval-augmented personalization method presented in LaMP (Salemi et al., 2023), where the user’s query is augmented with top $k$ retrieved items from the corresponding user’s history corpus. We take $k$=1, 2, 4 in this work.
  • Profile-Augmented Personalization (PAG): This method is taken from Richardson et al. (2023), in which the user’s input sequence would concatenate the user’s profile summarizing the user’s preference and behavior patterns. In our experiments, we generate user profiles using the vicuna-7B (Chiang et al., 2023) model. Moreover, the profile-augmented method could be combined with the retrieval augmentation. In this case, we take the number of retrieval items $k$=1 following the setting of Richardson et al. (2023).
For all baselines, we choose one of the most widely adopted open-source large language model Llama-2-7B (Touvron et al., 2023) as our base LLM and take BM25 (Trotman et al., 2014) for all retrieval operations to ensure efficient and fair comparison.
4.1.3. Evaluation Metrics
Following LaMP (Salemi et al., 2023), we use accuracy and F1-score for classification tasks (LaMP-1: personalized citation identification, LaMP-2N: personalized news categorization, and LaMP-2M: personalized movie tagging), MAE and RMSE for LaMP-3: personalized product rating, and adopt ROUGE-1 and ROUGE-L (Lin, 2004) for text generation tasks (LaMP-4: personalized news headline generation, LaMP-5: personalized scholarly title generation, LaMP-7: personalized tweet paraphrasing). Note that all metrics are the higher the better, except for RMSE and MAE used for the LaMP-3: personalized product rating task.
4.2. Main Results
Table 2 shows the performance on the test set of all seven public tasks in the LaMP benchmark. From the experimental results, we observe that
  • Impact of OPPU. Models equipped with OPPU outperform all corresponding personalization baseline methods across all seven tasks. Notably, in personalized classification tasks, OPPU achieves an average relative improvement of 17.38% and 8.89% on MAE and RMSE metrics on personalized product rating prediction task, as well as 11.87% accuracy and 7.56% F1-score performance gain on personalized movie tagging task. For personalized text generation tasks, we observe a 3.42% and 3.87% enhancement in ROUGE-1 and ROUGE-L scores, respectively, for personalized scholarly title generation.
  • Integrating non-parametric & parametric knowledge. Combining OPPU’s parametric knowledge stored in PEFT parameters and the non-parametric in retrieved items and user profiles, results in notable performance gains. For instance, averaging across all seven tasks, combining retrieval in OPPU will bring 1.93% and 2.48% relative improvement compared with the non-retrieval and non-OPPU yet retrieval version model, respectively. Moreover, integrating OPPU with user profiles would also bring 4.56% and 7.18% performance gain against non-profile and non-OPPU versions, respectively. Overall, integrating non-parametric knowledge from retrieval and user profile and parametric knowledge from personalized PEFT parameter in OPPU generally presents the best performance.
  • Impact of task & history format difference. In tasks like personalized citation identification, there’s a notable discrepancy between the format of user history and the task itself. Here, the user history comprises the user’s publication history, while the task involves a binary classification to identify the correct citation paper. This disparity is similarly observed in the personalized tweet paraphrasing task. In such cases, our proposed OPPU can robustly enhance performance. Specifically, in the personalized citation identification task, OPPU contributes to a notable increase of 3.48% in accuracy and 3.52% in the F1-score. This improvement is attributed to the provision of personalized context knowledge in a parametric manner via private PEFT.
  • Number of retrieved items. Our experimental results generally indicate that an increase in the number of retrieved items correlates with improved performance. However, we also observe that some data points don’t fit this trend, and we hypothesize that this inconsistency may arise from the retrieved items introducing noise and irrelevant behavior patterns, potentially complicating the model’s process of understanding user preferences.
Table 3. Performance under user behavior shift, where we remove the user behavior history highly similar to the query at hand. $k$ denotes the number of retrieved history items, and $k=0$ means non-retrieval. Armed with irrelevant user history, the retrieval-only method falls short and performs close to the non-personalized baseline, while OPPU shows stronger generalizability in the user behavior shift scenario.
LaMP Task History Type Non- Personalized Retrieval k=1 OPPU k=0 OPPU k=1
1 Acc F1 Acc F1 Acc F1 Acc F1
full .659 .657 .659 .657 .683 .682 .675 .674
irrelevant .626 .626 .683 .683 .699 .697
3 MAE RMSE MAE RMSE MAE RMSE MAE RMSE
full .223 .491 .214 .535 .179 .443 .196 .518
irrelevant .268 .583 .196 .463 .241 .559
5 R-1 R-L R-1 R-L R-1 R-L R-1 R-L
full .476 .415 .505 .445 .519 .442 .522 .457
irrelevant .475 .417 .493 .437 .490 .417
7 R-1 R-L R-1 R-L R-1 R-L R-1 R-L
full .527 .474 .571 .521 .539 .483 .579 .533
irrelevant .543 .495 .528 .482 .563 .523
4.3. Performance under User Behavior Shift
Recent studies have shown that retrieval-augmented generation methods tend to underperform when the retrieved corpus does not contain highly relevant documents (Petroni et al., 2020; Shi et al., 2023; Gupta et al., 2024). This issue often emerges in personalization contexts where the retrieval corpus is confined to a specific user’s behavior history and the user does not have a history of behavior closely matching their current queries. In our experiment, we simulate such a scenario where there is little or no overlap between the user’s behavior history corpus and their current query. Specifically, we employ the encoder-only language model DeBERTa-v3 (He et al., 2022) to extract features for the user’s historical behaviors and the query, then compute the cosine similarity between the query and all historical items to assess their relevance. By ranking the historical behaviors, we select the top 100 items with the lowest relevance scores as irrelevant user history.
As is demonstrated in Table 3, limiting user history to less relevant results in a marked decline in the performance of retrieval-based methods, often close to the performance of non-personalized approaches. Our proposed OPPU shows stronger robustness and generalization to these less relevant history behaviors and would even outperform using all user history items for private PEFT training. Moreover, the integration of both parametric and non-parametric knowledge (OPPU, k=1) exhibits enhanced robustness in personalized text generation tasks. In contrast, models utilizing only parametric user knowledge (OPPU, k=0) perform better in personalized text classification tasks.
Figure 3. Model performance on personalized movie tagging and personalized tweet paraphrasing for users with different numbers of behavior history.
Figure 4. Cosine similarities between personal PEFT parameters under personalized text classification and generation tasks.
4.4. Modeling Users with Different Active Levels
In the main experiment, we concentrate on modeling highly active users. However, it is worth noting that a significant number of users exhibit lower levels of activity, resulting in a comparatively short behavior history. To investigate the impact of user activity levels quantified by the number of historical behavioral items on model performance, we randomly chose 20 users from each range of active levels. As illustrated in Figure 3, equipped with OPPU, LLMs generally outperform the baseline methods across different user activity levels. Specifically, 1) the longer the user history length, the superiority of retrieval+OPPU over the baseline is generally larger. 2) The inclusion of non-parametric user knowledge via retrieval results in performance improvements compared to methods without retrieval. 3) Integrating the parametric knowledge in OPPU and non-parametric knowledge in retrieval generally shows the strongest performance over users across different active levels.
Figure 5. Performance of OPPU and retrieved-only baseline when the number of retrieved items $k$ increases.
4.5. Impact of Retrieved History Items
In this study, we alter the number of retrieved items of both retrieval-only baseline and retrieval+OPPU to gain a better understanding of the integration of non-parametric and parametric user knowledge. Figure 5 illustrates that as we increase the number of retrieved historical behavior items, both the retrieval-only baselines and the retrieval+OPPU approaches show improved performance. Interestingly, we observe that as the number of retrieved items $k$ becomes larger, the performance difference between the retrieval-only baseline and retrieval+OPPU narrows. This trend could be attributed to the longer logs of user behavior history in non-parametric prompts, which reduce the gap between the comprehensive user behavior history encapsulated in personalized PEFT parameters and the non-parametric user knowledge included in the prompts.
4.6. Similarities Between Personalized PEFTs
To gain a better understanding of how users’ behavior biases are encapsulated within their private PEFT parameters, we analyze the cosine similarities between these parameters across different users, as illustrated in Figure 4. Specifically, we select two representative tasks from text classification and generation categories respectively, then we compute the cosine similarities on the 100 users’ PEFT parameters in the test set. As shown in Figure 4, we observe that the private PEFT similarities generally range from 0.4 to 0.7. Interestingly, the personalized scholarly title generation task exhibits the highest average similarities, likely due to task-specific characteristics that entail less personal bias. Besides the absolute values of these similarities, the relative differences among various users provide additional insights. In personalized text classification tasks, the similarities tend to exhibit more variance, suggesting that some users have higher similarities compared to others. In contrast, the similarities in personalized text generation tasks remain relatively uniform. This pattern leads us to speculate that personal preferences in text generation tasks are more challenging to categorize, making it harder to distinctly group users based on their preferences. On the other hand, preferences in text classification tasks appear to be more identifiable and classifiable.
Table 4. Performance of OPPU with different ablated versions of user history configurations. $k$ refers to the number of retrieved items, and $k=0$ denotes non-retrieval. The best score is in bold and the second best is underlined.
Task in LaMP History Retrieval k=1 OPPU k=0 OPPU k=1
2M w/ desc. w/ tag Acc F1 Acc F1 Acc F1
$\checkmark$ 0.530 0.488 0.486 0.437 0.624 0.539
$\checkmark$ 0.567 0.514 0.499 0.44 0.634 0.548
$\checkmark$ $\checkmark$ 0.587 0.512 0.600 0.493 0.626 0.531
5 w/ abs. w/ title R-1 R-L R-1 R-L R-1 R-L
$\checkmark$ 0.493 0.422 0.497 0.434 0.495 0.449
$\checkmark$ 0.475 0.425 0.489 0.430 0.492 0.429
$\checkmark$ $\checkmark$ 0.505 0.445 0.519 0.442 0.522 0.457
Figure 6. We present a case study on a specific query from a user and analyze the responses generated by a non-personalized model in the personalized movie tagging task, a retrieval-augmented personalization model, and our proposed OPPU. It is shown that the retrieval-augmented personalization method can be easily distracted by less relevant user behavior history. In contrast, our OPPU demonstrates a more effective and comprehensive ability to capture the user’s behavior patterns.
Figure 7. Performance of OPPU on personalized movie tagging and personalized scholarly title generation tasks when equipped with different PEFT methods. We find that a larger proportion of trainable parameters generally results in better personalization performance.
4.7. Robustness against Task Formats
Our main results demonstrate that even with a history corpus that does not strictly follow the task format, our plug-in OPPU would bring significant performance improvement. In this study, we ablate the history format to test the robustness against user history format on personalized movie tagging (LaMP-2M) and personalized scholarly title generation (LaMP-5), from text classification and generation category respectively. Specifically, in both tasks, each user history item consists of both input and output aligned with the user query $x_{u}$ and output $y_{u}$ at hand. We ablate the history behavior items from the input and output side respectively and compare them with the retrieval baseline to test OPPU’s robustness against the history in mismatched format.
As is shown in Table 4, even with incomplete user behavior history that is not aligned with task format, OPPU still achieves relatively close performance with full history in text generation task. In the news categorization, LLM struggles to correctly classify with only parametric knowledge. However, integrating with retrieval augmentation, OPPU shows strong and robust performance, which can even outperform the full model tuned on the complete pairs of user history data. Overall, experimental results reveal that the integration of both non-parametric and parametric knowledge can make a robust model with different formats of user history.
4.8. On PEFT Method Choices
We envision the proposed OPPU as a versatile, PEFT-based LLM personalization framework compatible with various PEFT methods. This study demonstrates OPPU’s performance across different PEFT approaches. Beyond the most commonly adopted LoRA, we explore prompt tuning and $\mathrm{(IA)^{3}}$, which plug in external learnable parameters in the embedding space and scale the attention factor, respectively. As Figure 7 illustrates, the OPPU framework enhances performance with all three PEFT types, demonstrating the effectiveness and versatility of OPPU across various PEFT methods. Notably, LoRA typically delivers the highest performance, followed by $\mathrm{(IA)^{3}}$, and then prompt tuning. This hierarchy aligns with the proportion of trainable parameters in each method: LoRA at 0.01%, $\mathrm{(IA)^{3}}$ at 0.06%, and prompt tuning at 0.001%. These results suggest that a greater number of trainable parameters in a personalized PEFT method generally leads to improved personalization performance.
4.9. Case Study
To provide a qualitative understanding of OPPU, we conduct a case study focusing on an individual user in the personalized movie tagging task. As shown in Figure 6, the non-personalized method overlooks the user’s behavior history and bases its prediction solely on the user’s query input, leading to an evidently incorrect answer. The retrieval-based method incorporates user behavior history by retrieving the most relevant user history items, but it falls short of identifying the user’s most pertinent historical behaviors that mirror the query, leading to incorrect LLM output. We argue that retrieval augmentation only involves user history data via several retrieval examples, limiting the comprehensive understanding of user preferences. In contrast, our proposed OPPU employs a personalized PEFT module to effectively capture the user’s behavior preferences and patterns across the entire user history corpus. In this case, based on the user’s most frequently tags movies ”based on a book” in behavior history, OPPU successfully identifies this pattern and consequently provides the correct response.
5. Conclusion
LLMs personalization has emerged as a rapidly evolving research area, aiming to tailor LLMs’ emergent abilities to users’ unique needs. In this work, we introduce OPPU, which utilizes personalized PEFT parameters as a proxy for LLM personalization, demonstrating the advantage of model ownership and enhanced generalization under user behavior shift. By tuning these parameters with a user’s behavioral history, OPPU encapsulates the user’s behavioral history and patterns in private PEFT parameters. Furthermore, OPPU can be integrated with non-parametric user knowledge via history retrieval and user profile augmentation, exhibiting state-of-the-art performance across all seven public tasks in the LaMP benchmark. Additional experiments highlight OPPU’s versatility, robustness, and superiority in modeling users across different active levels. Our proposed OPPU framework paves the way for new opportunities in PEFT-based LLM personalization, enhancing the modularity of LLMs for more effective and democratizing personalization.
Appendix A Limitations
We identify two key limitations in OPPU. Firstly, limited by the dataset, we mainly focus on one specific task per user rather than examining user behaviors across multiple tasks and domains. For example, in the movie tagging task, users are solely engaged in that specific activity, without the inclusion of behaviors from other areas. Despite this, the OPPU framework is inherently adaptable to any text sequence generation task and is capable of conducting diverse user instructions across different tasks and domains. The exploration of LLM personalization across a broader range of tasks and domains remains an area for future investigation. Secondly, OPPU serves as a general framework that incorporates the entirety of a user’s behavior history into their private PEFT module. However, user interests are dynamic and may display inconsistencies or conflicts over time. Future research directions include examining methodologies for selecting the most relevant or valuable items from a user’s history and devising strategies to effectively manage any discrepancies or conflicts within this historical data.
Table 5. Hyperparameter settings of OPPU accross various tasks on LaMP benchmark. We find our hyperparameter settings robust across all 7 tasks.
Tasks rank #epoch lr R2 reg. batch size
LaMP-1: Personalized Citation Identification 8 3 $1e^{-5}$ $1e^{-2}$ 16
LaMP-2: Personalized News Categorization 8 3 $1e^{-5}$ $1e^{-2}$ 16
LaMP-2: Personalized Movie Tagging 8 3 $1e^{-5}$ $1e^{-2}$ 4
LaMP-3: Personalized Product Rating 8 3 $1e^{-5}$ $1e^{-2}$ 3
LaMP-4: Personalized News Headline Generation 8 2 $1e^{-5}$ $1e^{-1}$ 8
LaMP-5: Personalized Scholarly Title Generation 8 2 $1e^{-5}$ $1e^{-1}$ 4
LaMP-7: Personalized Tweet Paraphrasing 8 2 $1e^{-5}$ $1e^{-1}$ 8
Appendix B Ethics Statement
Privacy
Personalization in LLMs involves tailoring responses based on user-specific data, which may include sensitive or private information. The capacity of an LLM to adapt its outputs to individual users raises privacy concerns, as it might inadvertently reveal personal details. This underscores the importance of implementing robust privacy safeguards in LLM personalization, ensuring that personal data is handled respectfully and securely to prevent any unintended disclosures.
Data Bias
Personalizing LLMs heavily relies on the personal data fed into the system. If this personal data is biased or unrepresentative, the model’s outputs could potentially perpetuate these biases, leading to unfair or prejudiced responses. It is crucial to monitor and mitigate such biases in the personal data and the personalized model we obtain to ensure that personalized LLMs are fair and harmless in their responses.
Accessibility
By advancing the field of LLM personalization, we aim to enrich user interactions with AI systems. However, the complexity and resource-intensive nature of LLMs might pose accessibility challenges. Smaller entities or individual researchers with limited computational power and budgetary constraints might find it difficult to engage with advanced personalized LLMs, potentially widening the gap in AI research and application. It is essential to develop strategies that make personalized LLM technologies more accessible to a broader range of users and researchers, ensuring equitable progress in this domain.
Appendix C Hyperparameters
The hyperparameters of OPPU are presented in Table 5 to facilitate further research.
Figure 8. Efficiency analysis of OPPU, in which we alter the number of history items and average token per history item and record the training time.
Appendix D Efficiency Analysis
Personalization is a technique that aims at universally benefiting everyone, where scalability and efficiency are crucial factors in large-scale deployment. In this experiment, we study the training efficiency of our proposed OPPU. We specifically examine two critical factors: the number of user history items and the average token numbers per history item across classification and generation tasks. Given that the training of each user’s private PEFT can occur simultaneously or in a distributed manner, we choose not to consider the user count factor in this scenario, concentrating instead on the efficiency of training for an individual user. Initially, we set a consistent count of 100 whitespace-separated tokens for each history entry and vary the number of history items from 10 to 100. We then fix the history item count at 10 and adjust the token count from 10 to 100. The training time for each configuration, necessary for users to develop their personal PEFT modules. Presented in Figure 8, the results suggest that training time increases linearly with the number of user history items. Theoretically, training time grows quadratically with the increase in average tokens per history entry, yet our observations indicate a trend more akin to linear growth. It’s noteworthy that the longer training durations for personalized movie tagging tasks, as opposed to personalized tweet paraphrasing, are attributed to different training epochs.
Appendix E Scientific Artifacts
OPPU is built with the help of many existing scientific artifacts, including PyTorch (Paszke et al., 2019), Numpy (Harris et al., 2020), huggingface transformers (Wolf et al., 2020), and bitsandbytes (Dettmers et al., 2022). We will make the OPPU implementation publicly available to facilitate further research.
Appendix F Computation Resources Details
All experiments are implemented on a server with 3 NVIDIA A6000 GPU and Intel(R) Xeon(R) Silver 4210R CPU @ 2.40GHz with 20 CPU cores. Training 100 personal PEFT sequentially took around 12 minutes to 12 hours depending on the size of the behavior history corpus and the sequence length per history item.
Appendix G PEFT Cosine Similarity Details
Each user’s private PEFT parameters contain multiple learnable tensors, we first flatten the tensors and calculate the cosine similarities between corresponding private PEFT parameters, then average cosine similarities for each pair of PEFT modules. A pseudo-code using PyTorch is as follows:
def cosine_similarity(PEFT_1, PEFT_2):
    similarity_sum = 0
    count = 0
    for key in PEFT_1:
        if key in PEFT_2:
            v1 = PEFT_1[key].flatten()
            v2 = PEFT_2[key].flatten()
            dot = torch.dot(v1, v2)
            norm_1 = torch.linalg.norm(v1)
            norm_2 = torch.linalg.norm(v2)
            similarity = dot / (norm_1 * norm_2)
            similarity_sum += similarity
            count += 1
    return similarity_sum / count
Appendix H Task Details
We present the task details as follows to help readers gain a better understanding of the task format.
  • Personalized Citation Identification is a binary text classification task. Specifically, given user $u$ writes a paper $x$, the task aims to make the model determine which of the two candidate papers $u$ will cite in paper $x$ based on the user’s history data, which contains the publications of user $u$.
  • Personalized News Categorization is a 15-way text classification task to classify news articles written by a user $u$. Formally, given a news article $x$ written by user $u$, the language model is required to predict its category from the set of categories based on the user’s history data, which contains the user’s past article and corresponding category.
  • Personalized Movie Tagging is a 15-way text classification task to make tag assignments aligned with the user’s history tagging preference. Specifically, given a movie description $x$, the model needs to predict one of the tags for the movie $x$ based on the user’s historical movie-tag pairs.
  • Personalized Product Rating is a 5-way text classification task and can also be understood as a regression task. Given the user $u$’s historical review and rating pairs and the input review $x$, the model needs to predict the rating corresponding to $x$ selected from 1 to 5 in integer.
  • Personalized News Headline Generation is a text generation task to test the model’s ability to capture the stylistic patterns in personal data. Given a query $x$ that requests to generate a news headline for an article, as well as the user profile that contains the author’s historical article-title pairs, the model is required to generate a news headline specifically for the given user.
  • Personalized Scholarly Title Generation is a text generation task to test personalized text generation tasks in different domains. In this task, we require language models to generate titles for an input article $x$, given a user profile of historical article-title pairs for an author.
  • Personalized Tweet Paraphrasing is also a text generation task that tests the model’s capabilities in capturing the stylistic patterns of authors. Given a user input text $x$ and the user profile of historical tweets, the model is required to paraphrase $x$ into $y$ that follows the given user’s tweet pattern.
Appendix I Prompt Details
We present the prompt used in our experiments in this section, where the text in {BRACES} can be replaced with content specific to different users and queries.
I.1. Personalized Citation Identification
{USER PROFILE}
{RETRIEVED HISTORY}
Identify the most relevant reference for the listed publication by the researcher. Select the reference paper that is most closely related to the researcherś work. Please respond with only the number that corresponds to the reference.
paper title: {QUERY PAPER TITLE}
reference: [1] - {OPTION1} [2] - {OPTION2}
answer:
I.2. Personalized News Categorization
{USER PROFILE}
{RETRIEVED HISTORY}
Which category does this article relate to among the following categories? Just answer with the category name without further explanation. categories: [travel, education, parents, style & beauty, entertainment, food & drink, science & technology, business, sports, healthy living, women, politics, crime, culture & arts, religion]
article: {QUERY ARTICLE} category:
I.3. Personalized Movie Tagging
{USER PROFILE}
{RETRIEVED HISTORY}
Which tag does this movie relate to among the following tags? Just answer with the tag name without further explanation. tags: [sci-fi, based on a book, comedy, action, twist ending, dystopia, dark comedy, classic, psychology, fantasy, romance, thought-provoking, social commentary, violence, true story]
description: {QUERY DESCRIPTION} tag:
I.4. Personalized Product Rating
{USER PROFILE}
{RETRIEVED HISTORY}
What is the score of the following review on a scale of 1 to 5? just answer with 1, 2, 3, 4, or 5 without further explanation.
review: {QUERY REVIEW} score:
I.5. Personalized News Headline Generation
{USER PROFILE}
{RETRIEVED HISTORY}
Generate a headline for the following article.
article: {QUERY ARTICLE} headline:
I.6. Personalized Scholarly Title Generation
{USER PROFILE}
{RETRIEVED HISTORY}
Generate a title for the following abstract of a paper.
abstract: {QUERY ABSTRACT} title:
I.7. Personalized Tweet Paraphrasing
{USER PROFILE}
{RETRIEVED HISTORY}
Following the given pattern, paraphrase the following text into tweet without any explanation before or after it.
text: {QUERY TEXT} tweet: